### v1

# Holographic Aspects of Non-minimal $R^{3}F^{(a)}_{\mu\alpha}F^{(a)\mu\alpha}$ Black Brane

Abstract
This work investigates a modified theory of gravity where the Einstein-Hilbert action, including a cosmological constant, is non-minimally coupled to a Yang-Mills field via an $R^{3}F^{(a)}_{\mu\alpha}F^{(a)\mu\alpha}$ interaction term. We derive a black brane solution for this model, accurate to the first order in the coupling parameter. Using gauge/gravity duality techniques, we then compute two key holographic transport coefficients: the color non-abelian direct current (DC) conductivity and the ratio of shear viscosity to entropy density. Our analysis reveals that both transport coefficients are modified by the non-minimal coupling, with the conductivity bound violated for positive $q_{2}$ and the Kovtun-Son-Starinets (KSS) bound for shear viscosity violated for negative $q_{2}$. In the limit where the non-minimal coupling vanishes, our results consistently reduce to those of the standard Yang-Mills Schwarzschild Anti-de Sitter (AdS) black brane.
PACS numbers: 11.10.Jj, 11.10.Wx, 11.15.Pg, 11.25.Tq
Keywords: AdS/CFT duality, DC Conductivity, Black brane, Shear viscosity, Entropy density
1 Introduction
Black holes, as solutions to Einstein’s field equations characterized by their event horizons, represent one of the most intriguing predictions of general relativity. The classical notion that nothing escapes the event horizon was revolutionized by the discovery of Hawking radiation, providing a quantum mechanical window into their thermodynamic properties [1]. Understanding the quantum and thermodynamic aspects of black holes remains a central pursuit in theoretical physics.
A profound development in this field is the Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence [2]-[5]. This duality posits an equivalence between a gravitational theory in an asymptotically AdS spacetime and a conformal field theory (CFT) residing on its boundary. This framework provides a powerful tool for analyzing strongly coupled quantum field theories by mapping them to a classical gravity problem. In particular, the long-wavelength, low-frequency limit of this duality leads to the fluid-gravity correspondence [6]-[11], where the boundary theory is described by hydrodynamics. Within this framework, transport coefficients of the boundary fluid, such as conductivity and shear viscosity, can be computed from the gravitational dual using methods like the Green-Kubo formula [6].
A compelling direction in gravitational model-building involves non-minimal couplings between curvature invariants and matter fields. Such couplings, which modify the standard Einstein-Hilbert action, can offer alternative explanations for cosmological phenomena like inflation and late-time acceleration [12]. Among the various possibilities, coupling the Ricci scalar to the Yang-Mills field strength invariant, $F^{(a)}_{\mu\alpha}F^{(a)\mu\alpha}$, represents a particularly rich class of non-minimal models [13, 14]. These theories have been shown to admit exact solutions for stars [15, 16], wormholes [17, 18], and black holes [19, 20], often utilizing the Wu-Yang ansatz. While previous studies have explored lower-order couplings like $RF^{2}$ [21] and $R_{\mu\alpha\nu\beta}F^{(a)\mu\alpha}F^{(a)\nu\beta}$ [22] the present work extends this program to the next order by investigating the holographic implications of an $R^{3}F^{2}$ coupling.
In holographic systems, transport coefficients are often conjectured to obey universal bounds. A famous example is the Kovtun-Son-Starinets (KSS) bound, which posits a lower limit for the ratio of shear viscosity to entropy density, $\frac{\eta}{s}\geq\frac{1}{4\pi}$ [23]. While this bound is saturated in Einstein-Hilbert gravity, it can be violated in theories with higher-derivative terms [23, 24], massive gravity [25]-[26], and various scalar-tensor theories [27, 28, 29]. Similarly, a proposed bound for the DC conductivity, $\sigma\geq 1$, has been found to be violated in several holographic setups, including massive gravity [30], theories with background fields [31], nonlinear model [32], non-abelian Born-Infeld models [33, 34], and Non-abelian exponential Yang-Mills AdS black brane [35].
The primary objective of this paper is to construct a black brane solution in a non-minimal Einstein-Yang-Mills theory featuring a specific $R^{3}F^{2}$ coupling and to study its holographic characteristics. We will calculate the color non-abelian DC conductivity and the shear viscosity to entropy density ratio for this system. A central question we aim to address is whether the presence of this higher-order non-minimal coupling preserves or violates the aforementioned universal bounds for transport coefficients.
2 Non-minimal $R^{3}F^{(a)}_{\mu\alpha}F^{(a)\mu\alpha}$ AdS Black Brane Solution
The non-minimal Einstein-Yang-Mills theory with a negative cosmological constant can be described in terms of the action functional below[13],[36],[37],
$S=\int d^{4}x\sqrt{-g}\bigg[\frac{1}{\kappa}(R-2\Lambda)-\frac{q_{1}}{4}F^{(a)}_{\mu\alpha}F^{(a)\mu\alpha}+q_{2}R^{3}F^{(a)}_{\mu\alpha}F^{(a)\mu\alpha}\bigg],$ (1)
where $\kappa$ is the gravitational constant, $R$ is the Ricci scalar, $\Lambda=-\frac{3}{L^{2}}$ is the cosmological constant, $L$ is the AdS radius, $q_{1}$ is a constant dimensionless parameter, and $\mathcal{F}=\text{Tr}(F_{\mu\nu}^{(a)}F^{(a)\mu\nu})$ is the Yang-Mills invariant. The dimensionful coupling constant $q_{2}$ (with dimensions $L]^{6}$) characterizes the non-minimal interaction between the gauge field and curvature.
The Yang-Mills field tensor is given by:
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}-i[A_{\mu},A_{\nu}],$ (2)
where the gauge coupling constant is set to unity, and $A_{\nu}$ represents the Yang-Mills potentials.
Variation of the action (1) with respect to the spacetime metric $g_{\mu\nu}$ yields the field equations:
$R_{\mu\nu}-\tfrac{1}{2}g_{\mu\nu}R+\Lambda g_{\mu\nu}=\kappa T^{\text{(eff)}}_{\mu\nu}$ (3)
where,
$T^{\text{(eff)}}_{\mu\nu}=q_{1}T^{\text{(YM)}}_{\mu\nu}+q_{2}T^{(I)}_{\mu\nu}$ (4)
$T^{\text{(YM)}}_{\mu\nu}=-\tfrac{1}{8}g_{\mu\nu}F^{(a)}_{\alpha\beta}F^{{}^{(a)}\alpha\beta}+\tfrac{1}{2}F_{\mu}{}^{(a)\alpha}F^{(a)}_{\nu\alpha}$ (5)
$T^{(I)}_{\mu\nu}=-3F^{(a)}_{\alpha\beta}F^{(a)\,\alpha\beta}R_{\mu\nu}R^{2}-2F_{\mu}^{(a)\,\alpha}F^{(a)}_{\nu\alpha}R^{3}$
$+\tfrac{1}{2}F^{(a)}_{\alpha\beta}F^{(a)\,\alpha\beta}g_{\mu\nu}R^{3}-24F^{(a)\,\beta\gamma}g_{\mu\nu}R\nabla_{\alpha}F^{(a)}_{\beta\gamma}\nabla^{\alpha}R\$
$-6F^{(a)}_{\beta\gamma}F^{(a)\,\beta\gamma}g_{\mu\nu}\nabla_{\alpha}R\nabla^{\alpha}R-6F^{(a)\alpha\beta}g_{\mu\nu}R^{2}\nabla_{\gamma}\nabla^{\gamma}F^{(a)}_{\alpha\beta}$
$-6F^{(a)}_{\alpha\beta}F^{(a)\,\alpha\beta}g_{\mu\nu}R\nabla_{\gamma}\nabla^{\gamma}R-6g_{\mu\nu}R^{2}\nabla_{\gamma}F^{(a)\,\alpha\beta}\nabla^{\gamma}F^{(a)}_{\alpha\beta}$
$+3F^{(a)\,\alpha\beta}R^{2}\nabla_{\mu}\nabla_{\nu}F^{(a)}_{\alpha\beta}+6R^{2}\nabla_{\mu}F^{(a)\,\,\alpha\beta}\ \nabla_{\nu}F^{(a)}_{\alpha\beta}$
$+12F^{(a)\alpha\beta}R\nabla_{\mu}R\nabla_{\nu}F^{(a)}_{\alpha\beta}+12F^{(a)\,\alpha\beta}R\nabla_{\mu}F^{(a)}_{\alpha\beta}\nabla_{\nu}R$
$+6F^{(a)}_{\alpha\beta}F^{(a)\,\alpha\beta}\nabla_{\mu}R\nabla_{\nu}R+3F^{(a)\,\alpha\beta}R^{2}\nabla_{\nu}\nabla_{\mu}F^{(a)}_{\alpha\beta}$
$+6F^{(a)}_{\alpha\beta}F^{(a)\,\alpha\beta}R\nabla_{\nu}\nabla_{\mu}R.$ (6)
Variation with respect to $A_{\mu}$ yields the gauge field equations:
$\nabla_{\mu}\Big(-\frac{1}{2}q_{1}F^{(a)\mu\nu}+2q_{2}F^{(a)\mu\nu}R^{3}\Big)=0.$ (7)
For our four-dimensional spacetime with two-dimensional spatial symmetry, we employ the metric ansatz:
$ds^{2}=-e^{-2H(r)}f(r)dt^{2}+\frac{dr^{2}}{f(r)}+\frac{r^{2}}{L^{2}}(dx^{2}+dy^{2}).$ (8)
where the gauge group corresponds to the diagonal generator of the Cartan subalgebra of $SU(2)$ [40].
The field equations derived from Eqs. (7-9) yield the tttt-component:
We apply the ansatz 8 for solving Eq.(7) where the potential 1-form is expressed by,
${\bf{A}}^{(a)}=\frac{i}{2}h(r)dt\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$, (9)
The gauge group is the diagonal generator of the Cartan subalgebra of $SU(2)$ [40].
Now, we can write the field equations of motion by consideration of Eqs.(7-9).
The $tt$-component of the field equations of motion is as follows,
$r^{4}\left(4rf_{0}^{\prime}(r)+4f_{0}(r)+\kappa q_{1}r^{2}e^{2H_{0}(r)}h_{0}^{\prime}(r)^{2}+4\Lambda r^{2}\right)+q_{2}(...)=0,$ (10)
for $q_{2}=0$ we have,
$r^{4}\left(4rf_{0}^{\prime}(r)+4f_{0}(r)+\kappa q_{1}r^{2}e^{2H_{0}(r)}h_{0}^{\prime}(r)^{2}+4\Lambda r^{2}\right)=0,$ (11)
The $rr$-component of the field equations of motion is as follows,
$r^{4}\left(4rf_{0}^{\prime}(r)-8rf_{0}(r)H_{0}^{\prime}(r)+4f_{0}(r)+\kappa q_{1}r^{2}e^{2H_{0}(r)}h_{0}^{\prime}(r)^{2}+4\Lambda r^{2}\right)+q_{2}(...)=0,$ (12)
in which $rr$- part of Einstein equations for $q_{2}=0$ should be the same as Eq.(11). Therefore, we conclude that $H_{0}(r)=0$.
Non-zero component of Eq.(7) is as follows,
$B_{1}(r)h^{\prime\prime}(r)+B_{2}(r)h^{\prime}(r)=0,$ (13)
where,
$B_{1}(r)=2q_{1}f(r)r^{7}-8q_{2}f(r)rA(r)^{3},$ (14)
$B_{2}(r)=4q_{1}f(r)r^{6}+2q_{1}f(r)r^{7}H^{\prime}(r)+8q_{2}f(r)(4-rH^{\prime}(r))A(r)^{3}$
[…]
with the auxiliary functions:
$A(r)=rf^{\prime}(r)(-4+3rH^{\prime}(r))-r^{2}f^{\prime\prime}(r)+2f(r)\left[-(1-rH^{\prime}(r))^{2}+r^{2}H^{\prime\prime}(r)\right],$ (15)
$C(r)=-4f(r)H^{\prime}(r)+f^{\prime}(r)\left[6-10rH^{\prime}(r)+2r^{2}H^{\prime}(r)^{2}-5r^{2}H^{\prime\prime}(r)\right]$ (16)
$+r\left[(6-3rH^{\prime}(r))f^{\prime\prime}(r)+rf^{\prime\prime\prime}(r)+f(r)\left(-8H^{\prime\prime}(r)+4H^{\prime}(r)^{2}+4rH^{\prime}(r)H^{\prime\prime}(r)-2rH^{\prime\prime\prime}(r)\right)\right].$
The solution of $h(r)$ is as,
$h(r)=-C_{1}\int^{r}\frac{e^{-H(u)}u^{4}}{q_{1}u^{6}+4q_{2}B(u)^{3}}du+C_{2},$ (17)
$B(u)=-3u^{2}f^{\prime}(u)H^{\prime}(u)+u^{2}f^{\prime\prime}(u)+4uf^{\prime}(u)+2u^{2}f(u)H^{\prime}(u)^{2}$
$-2u^{2}f(u)H^{\prime\prime}(u)-4uf(u)H^{\prime}(u)+2f(u).$ (18)
As these equations lack analytical solutions, we employ a perturbative approach to first order in $q_{2}$ [41, 42]. We expand:
$f(r)=f_{0}(r)+q_{2}f_{1}(r),$ (19)
$h(r)=h_{0}(r)+q_{2}h_{1}(r),$ (20)
$H(r)=H_{0}(r)+q_{2}H_{1}(r),$ (21)
where $f_{0}(r)$, $h_{0}(r)$ and $H_{0}(r)$ are the leading order solutions of Einstein-Yang-Mills AdS black brane in four dimensions.
The zeroth-order solutions are:
$h_{0}(r)=C_{2}-C_{1}\int^{r}\frac{1}{q_{1}u^{2}}du=Q(\frac{1}{r}-\frac{1}{r_{h}}),$ (22)
where $C_{1}=q_{1}Q$ and $C_{2}=-\frac{Q}{r_{h}}$.
$f_{0}(r)=-\frac{2m_{0}}{r}-\frac{\Lambda r^{2}}{3}-\frac{\kappa q_{1}}{4r}\int^{r}u^{2}h_{0}^{\prime 2}du=-\frac{2m_{0}}{r}-\frac{\Lambda r^{2}}{3}+\frac{\kappa q_{1}Q^{2}}{4r^{2}},$ (23)
$H_{0}(r)=0.$ (24)
[…]
$2m_{0}=-\frac{\Lambda r_{h}^{3}}{3}+\frac{\kappa q_{1}Q^{2}}{4r_{h}}.$ (25)
By plugging Eq.(25) in Eq.(23) we have,
$f_{0}(r)=\frac{\Lambda}{3r}(r_{h}^{3}-r^{3})+\frac{\kappa q_{1}Q^{2}}{4r}(\frac{1}{r}-\frac{1}{r_{h}}).$ (26)
Eq.(10) and Eq.(12) should be the same up to first order of $q_{2}$. Therefore, $H_{1}$ is calculated as follows,
$H_{1}(r)=C_{1}+\frac{3\kappa}{r^{4}}h_{0}^{\prime}(r)\left(2f_{0}(r)+4rf_{0}^{\prime}(r)+r^{2}f_{0}^{\prime\prime}(r)\right)$
$\left[2r\left(2f_{0}(r)+4rf_{0}^{\prime}(r)+r^{2}f_{0}^{\prime\prime}(r)\right)h_{0}^{\prime\prime}(r)+h_{0}^{\prime}(r)\left(-10f_{0}(r)-8rf_{0}^{\prime}(r)+7r^{2}f_{0}^{\prime\prime}(r)+2r^{3}f_{0}^{\prime\prime\prime}(r)\right)\right]$
So we have,
$H_{1}(r)=-\frac{240\kappa Q^{2}\Lambda^{2}}{r^{4}}+C_{3},$ (28)
where $C_{3}$ is a dimensionless integration constant. Since the metric of our model is asymptotic AdS near the boundary $r\to\infty$ [[43]], so $C_{3}=0$. It also guarantees the speed of light to unity on the theory of boundary.
$h_{1}(r)=C_{4}+C_{5}\int^{r}\frac{q_{1}u^{6}H_{1}+4\left(2f_{0}+u\left(4f_{0}^{\prime}+uf_{0}^{\prime\prime}\right)\right)^{3}}{q_{1}^{2}u^{8}}du=C_{4}+\frac{48\kappa C_{5}Q^{2}\Lambda^{2}}{q_{1}r^{5}}+\frac{256C_{5}\Lambda^{3}}{q_{1}^{2}r},$ (29)
where $C_{5}=q_{1}^{2}Q$. $C_{4}$ is determined by applying this condition $h_{1}(r_{h})=0$.
$h_{1}(r)=48q_{1}\kappa Q^{3}\Lambda^{2}(\frac{1}{r^{5}}-\frac{1}{r_{h}^{5}})+256Q\Lambda^{3}(\frac{1}{r}-\frac{1}{r_{h}}),$ (30)
by substituting $H_{0}(r)=0$ we have,
$f_{1}(r)=\frac{1}{r}\int^{r}\frac{1}{2u^{4}}D(u)du+\frac{C_{6}}{r},$ (31)
where $D(u)$ is given by:
$D(u)=16\kappa f_{0}(u)^{3}h_{0}^{\prime}(u)\left(-25h_{0}^{\prime}(u)+12uh_{0}^{\prime\prime}(u)\right)$
$+48\kappa f_{0}(u)^{2}uh_{0}^{\prime}(u)\left(f_{0}^{\prime}(u)\left(-25h_{0}^{\prime}(u)+17uh_{0}^{\prime\prime}(u)\right)\right$.
$\quad\left.+u\left(f_{0}^{\prime\prime}(u)\left(3h_{0}^{\prime}(u)+4uh_{0}^{\prime\prime}(u)\right)+2uh_{0}^{\prime}(u)f_{0}^{(3)}(u)\right)\right)$
$+\kappa u^{3}h_{0}^{\prime}(u)\left(-q_{1}H_{1}(u)u^{3}h_{0}^{\prime}(u)-q_{1}u^{3}h_{1}^{\prime}(u)\right$.
$\quad+4\left(4f_{0}^{\prime}(u)+uf_{0}^{\prime\prime}(u)\right)\left(-2u^{2}h_{0}^{\prime}(u)f_{0}^{\prime\prime}(u)^{2}\right$.
$\quad+f_{0}^{\prime}(u)^{2}\left(-26h_{0}^{\prime}(u)+12uh_{0}^{\prime\prime}(u)\right)$
$\quad\left.\left.+uf_{0}^{\prime}(u)\left(-f_{0}^{\prime\prime}(u)\left(h_{0}^{\prime}(u)-3uh_{0}^{\prime\prime}(u)\right)+3uh_{0}^{\prime}(u)f_{0}^{(3)}(u)\right)\right)\right)$
$+4f_{0}(u)u^{2}\left(u^{3}H_{1}^{\prime}(u)\right$.
$\quad+3\kappa h_{0}^{\prime}(u)\left(f_{0}^{\prime}(u)^{2}\left(-84h_{0}^{\prime}(u)+80uh_{0}^{\prime\prime}(u)\right)\right$.
$\quad+u^{2}f_{0}^{\prime\prime}(u)\left(f_{0}^{\prime\prime}(u)\left(13h_{0}^{\prime}(u)+4uh_{0}^{\prime\prime}(u)\right)+4uh_{0}^{\prime}(u)f_{0}^{(3)}(u)\right)$
$\quad\left.\left.+2uf_{0}^{\prime}(u)\left(2f_{0}^{\prime\prime}(u)\left(10h_{0}^{\prime}(u)+9uh_{0}^{\prime\prime}(u)\right)+9uh_{0}^{\prime}(u)f_{0}^{(3)}(u)\right)\right)\right),$
$f_{1}(r)$ is calculated as,
$f_{1}(r)=$ $\frac{C_{6}}{r}+\frac{672m_{0}\kappa Q^{2}\Lambda^{2}}{r^{5}}-\frac{96q_{1}\kappa^{2}Q^{4}\Lambda^{2}}{r^{6}}+\frac{24q_{1}^{2}\kappa^{2}Q^{4}\Lambda^{2}}{r^{6}}$ (32)
$+\frac{128q_{1}\kappa Q^{2}\Lambda^{3}}{r^{2}}+\frac{96\kappa Q^{2}\Lambda^{3}}{r^{2}},$
where,
$C_{6}=-2m_{1}.$ (33)
Therefore, we have, by inserting Eq.(33) in Eq.(32) we have,
$f_{1}(r)=-\frac{2m_{1}}{r}+\frac{672m_{0}\kappa Q^{2}\Lambda^{2}}{r^{5}}-\frac{96q_{1}\kappa^{2}Q^{4}\Lambda^{2}}{r^{6}}+\frac{24q_{1}^{2}\kappa^{2}Q^{4}\Lambda^{2}}{r^{6}}$
$+\frac{128q_{1}\kappa Q^{2}\Lambda^{3}}{r^{2}}+\frac{96\kappa Q^{2}\Lambda^{3}}{r^{2}},$ (34)
Assembling all pieces, the complete perturbative solution to linear order in $q_{2}$ is:
$h(r)=Q(\frac{1}{r}-\frac{1}{r_{h}})+q_{2}\Big\{48q_{1}\kappa Q^{3}\Lambda^{2}(\frac{1}{r^{5}}-\frac{1}{r_{h}^{5}})+256Q\Lambda^{3}(\frac{1}{r}-\frac{1}{r_{h}}\Big\},$ (35)
$H(r)=-q_{2}\frac{240\kappa Q^{2}\Lambda^{2}}{r^{4}},$ (36)
$f(r)=-\frac{2m_{0}}{r}-\frac{\Lambda r^{2}}{3}+\frac{\kappa q_{1}Q^{2}}{4r^{2}}$
$+q_{2}\Big(\frac{672m_{0}\kappa Q^{2}\Lambda^{2}}{r^{5}}-\frac{96q_{1}\kappa^{2}Q^{4}\Lambda^{2}}{r^{6}}$
$+\frac{24q_{1}^{2}\kappa^{2}Q^{4}\Lambda^{2}}{r^{6}}+\frac{128q_{1}\kappa Q^{2}\Lambda^{3}}{r^{2}}+\frac{96\kappa Q^{2}\Lambda^{3}}{r^{2}}\Big).$ (37)
We perform a final reparametrization by defining $\bar{m}_{0}=m_{0}+q_{2}m_{1}$, which absorbs the $m_{1}$ term into the leading-order mass. The final form of the blackening function becomes:
$f(r)=-\frac{2\bar{m}_{0}}{r}-\frac{\Lambda r^{2}}{3}+\frac{\kappa q_{1}Q^{2}}{4r^{2}}$
$+q_{2}\Big(\frac{672m_{0}\kappa Q^{2}\Lambda^{2}}{r^{5}}-\frac{96q_{1}\kappa^{2}Q^{4}\Lambda^{2}}{r^{6}}$
$+\frac{24q_{1}^{2}\kappa^{2}Q^{4}\Lambda^{2}}{r^{6}}+\frac{128q_{1}\kappa Q^{2}\Lambda^{3}}{r^{2}}+\frac{96\kappa Q^{2}\Lambda^{3}}{r^{2}}\Big).$ (38)
This constitutes the complete perturbative solution to linear order in $q_{2}$.
We aim to calculate the non-abelian DC conductivity and the ratio of shear viscosity to entropy density as two important transport coefficients using fluid-gravity duality to describe the holographic dual of our model.
3 Color DC Conductivity
We employ the Green-Kubo formula [8] for calculating the non-abelian color DC conductivity:
$\sigma^{ij}(k_{\mu})=-\lim_{\omega\to 0}\frac{1}{\omega}\Im G^{ij}(k_{\mu}).$ (39)
The retarded Green’s function is computed using AdS/CFT duality. First, we perturb the gauge field as $A_{\mu}\to A_{\mu}+\tilde{A}_{\mu}$ and substitute it into the action. We expand the resulting action (1) to second order in the perturbation. The Green’s function is then obtained by taking the second derivative with respect to the boundary value of the gauge field [8]:
$\sigma^{\mu\nu}(\omega)=\frac{1}{i\omega}\langle J^{\mu}(\omega)J^{\nu}(-\omega)\rangle=\frac{\delta^{2}S}{\delta\tilde{A}^{0}_{\mu}\delta\tilde{A}^{0}_{\nu}},$ (40)
where $\tilde{A}^{0}_{\nu}$ represents the gauge field perturbation at the boundary. The boundary exhibits $SO(2)$ symmetry, which ensures that the conductivity is a scalar quantity:
$\sigma^{ij}_{ab}=\sigma_{ab}\delta^{ij}.$ (41)
We consider the perturbation of the gauge field as $\tilde{A}_{x}=\tilde{A}_{x}(r)e^{-i\omega t}$, where $\omega$ is small due to the hydrodynamic regime.
Substituting the perturbation into the action Eq. (1) and keeping terms up to second order in $\tilde{A}$ yields:
$S^{(2)}$ $=-\int d^{4}x\frac{2e^{-H}}{fr^{6}}\Bigg[-f^{2}\left((\partial_{r}\tilde{A}_{x}^{(1)})^{2}+(\partial_{r}\tilde{A}_{x}^{(2)})^{2}+(\partial_{r}\tilde{A}_{x}^{(3)})^{2}\right)$
$+e^{2H}\left((\tilde{A}_{x}^{(1)})^{2}+(\tilde{A}_{x}^{(2)})^{2}\right)(\omega^{2}+h^{2})+e^{2H}\omega^{2}(\tilde{A}_{x}^{(3)})^{2}\Bigg]\left(q_{1}r^{6}+4q_{2}\Psi^{3}\right),$ (42)
where:
$\mathcal{Y}=q_{1}r^{6}+4q_{2}\Psi^{3},$ (43)
with:
$\Psi=r\left(f^{\prime}(4-3rH^{\prime})+rf^{\prime\prime}\right)+2f\left((-1+rH^{\prime})^{2}-r^{2}H^{\prime\prime}\right).$ (44)
Varying the action $S^{(2)}$ with respect to $\tilde{A}_{x}^{(1)}$ gives:
$-\frac{re^{2H(r)}\tilde{A}_{x}^{(1)}(\omega^{2}+h(r)^{2})}{f(r)}\mathcal{Y}+\left(6f(r)-rf^{\prime}(r)+rf(r)H^{\prime}(r)\right)\tilde{A}_{x}^{(1)^{\prime}}\mathcal{Y}$
$-rf(r)\left(\tilde{A}_{x}^{(1)^{\prime}}\mathcal{Y}\right)^{\prime}=0.$ (45)
The equation for $\tilde{A}_{x}^{(2)}$ is identical to that for $\tilde{A}_{x}^{(1)}$. The equation for $\tilde{A}_{x}^{(3)}$ is:
$r^{4}f^{\prime}(r)\tilde{A}_{x}^{(3)^{\prime}}\left(-q_{1}r^{3}+4q_{2}\left(f^{\prime}(r)(-4+3rH^{\prime}(r))-rf^{\prime\prime}(r)\right)^{3}\right)-\frac{re^{2H(r)}\omega^{2}\tilde{A}_{x}^{(3)}}{f(r)}\mathcal{Y}$ (46)
$+r^{4}f(r)\left[\tilde{A}_{x}^{(3)^{\prime\prime}}\left(-q_{1}r^{3}+4q_{2}\left(f^{\prime}(r)(-4+3rH^{\prime}(r))-rf^{\prime\prime}(r)\right)^{3}\right)+\tilde{A}_{x}^{(3)^{\prime}}\mathcal{Y}^{\prime}\right]$
[…]
$\left.+\tilde{A}_{x}^{(3)^{\prime}}\left(6+r\left(-2rH^{\prime}(r)^{2}+r^{2}H^{\prime}(r)^{3}-H^{\prime}(r)(5+7r^{2}H^{\prime\prime}(r))+3r(2H^{\prime\prime}(r)+rH^{\prime\prime\prime}(r))\right)\right)\right]$
[…]
$\left.\times\left(-(-1+rH^{\prime}(r))^{2}+r^{2}H^{\prime\prime}(r)\right)+\tilde{A}_{x}^{(3)^{\prime}}\mathcal{Y}^{\prime\prime}\right]$
[…]
$\left.\times\left(-(-1+rH^{\prime}(r))^{2}+r^{2}H^{\prime\prime}(r)\right)+\tilde{A}_{x}^{(3)^{\prime}}\mathcal{Y}^{\prime\prime\prime}\right]=0.$
Using the near-horizon behavior $f_{0}\sim 4\pi f_{0}^{\prime}(r_{h})(r-r_{h})$ and $f_{1}\sim 4\pi f_{1}^{\prime}(r_{h})(r-r_{h})$, we solve Eqs. (3) and (46) near the event horizon. Since $\tilde{A}_{x}^{(a)}$ must be ingoing at the horizon, we adopt the ansatz:
$\tilde{A}_{x}^{(a)}\sim(r-r_{h})^{z_{a}},\qquad a=1,2,3,$ (47)
where:
$z_{1}$ $=z_{2}=\pm i\frac{\sqrt{h(r_{h})^{2}+\omega^{2}}}{4\pi T},$ (48)
$z_{3}$ $=\pm i\frac{\omega}{4\pi T}.$ (49)
The Hawking temperature of the black brane is:
$T=\frac{1}{2\pi}\left[\frac{1}{\sqrt{g_{rr}}}\frac{d}{dr}\sqrt{-g_{tt}}\right]\Bigg|_{r=r_{h}}=\frac{e^{-H(r_{h})}f^{\prime}(r_{h})}{4\pi}.$ (50)
To solve for $\tilde{A}_{x}^{(a)}$ from the horizon to the boundary, we use the ansatz:
$\tilde{A}_{x}^{(1)}$ $=\tilde{A}^{(1)}_{\infty}\left(\frac{-3f}{\Lambda r^{2}}\right)^{z_{1}}\left(1+i\omega b_{1}(r)+\cdots\right),$ (51)
$\tilde{A}_{x}^{(2)}$ $=\tilde{A}^{(2)}_{\infty}\left(\frac{-3f}{\Lambda r^{2}}\right)^{z_{2}}\left(1+i\omega b_{2}(r)+\cdots\right),$ (52)
$\tilde{A}_{x}^{(3)}$ $=\tilde{A}^{(3)}_{\infty}\left(\frac{-3f}{\Lambda r^{2}}\right)^{z_{3}}\left(1+i\omega b_{3}(r)+\cdots\right),$ (53)
where $\tilde{A}^{(a)}_{\infty}$ represents the boundary values, and $z_{i}$ correspond to the negative signs in Eqs. (48) and (49), selecting the ingoing mode.
Substituting Eq. (53) into Eq. (46) and keeping terms to first order in $\omega$ yields:
$-r^{4}f^{\prime\prime}(r)\left(q_{1}+4q_{2}f^{\prime\prime}(r)^{3}\right)+\mathcal{B}_{4}(r)+32q_{2}E_{1}(r)^{2}E_{6}(r)f(r)^{4}$
$+16q_{2}rE_{1}(r)E_{7}(r)f(r)^{3}-24q_{2}r^{2}E_{9}(r)f(r)^{2}+r^{3}E_{8}(r)f(r)=0,$ (54)
where:
$E_{1}(r)$ $=(-1+rH^{\prime}(r))^{2}-r^{2}H^{\prime\prime}(r),$
$\mathcal{B}_{4}(r)$ $=-4q_{2}f^{\prime}(r)^{4}(4-3rH^{\prime}(r))^{2}E_{2}(r)-12q_{2}r^{2}f^{\prime}(r)^{2}f^{\prime\prime}(r)E_{3}(r)$
$\quad+4q_{2}rf^{\prime}(r)^{3}(-4+3rH^{\prime}(r))E_{4}(r)+r^{3}f^{\prime}(r)E_{5}(r).$ (55)
The auxiliary functions $E_{2}(r)$ through $E_{9}(r)$ contain the remaining polynomial structure in derivatives of $f$, $H$, and $b_{3}$ up to third order. Their complete definitions are provided in the appendix due to their length.
Imposing regularity of $\tilde{A}_{x}^{(3)}$ on the event horizon, the function $b_{3}(r)$ in Eq. (53) is determined as:
$b_{3}(r)=C_{8}+\int^{r}\frac{e^{H(s)}s^{6}}{f(s)\mathcal{Y}(s)}\left[C_{7}+\mathcal{I}_{1}(s)\right]ds,$ (56)
where:
$\mathcal{I}_{1}(s)$ $=\int^{s}e^{-H(t)}t^{-4}\mathcal{L}(t)dt,$ (57)
$\mathcal{Y}(s)$ $=q_{1}s^{6}+4q_{2}\Psi(s)^{3},$ (58)
and $C_{7}$ and $C_{8}$ are integration constants. The function $\mathcal{L}(t)$ is defined in the appendix.
The solution for $b_{3}(r)$ near the event horizon is:
$b_{3}\approx\left(C_{7}+\mathcal{I}_{1}(r_{h})\right)\log(r-r_{h})+\text{finite terms}.$ (59)
Regularity at the horizon requires:
$C_{7}=-\mathcal{I}_{1}(r_{h}).$ (60)
Using the solution for $\tilde{A}_{x}^{(3)}$ in Eq. (3) and varying with respect to $\tilde{A}^{(3)}_{\infty}$, the Green’s function is:
$G_{xx}^{(33)}(\omega,\vec{0})=-i\omega\frac{C_{7}}{q_{1}f^{\prime}(r_{h})}=\frac{i\omega}{q_{1}f^{\prime}(r_{h})}\mathcal{I}_{1}(r_{h}).$ (61)
The conductivity is then:
$\sigma_{xx}^{(33)}=-\frac{\mathcal{I}_{1}(r_{h})}{q_{1}f^{\prime}(r_{h})}=1-\frac{q_{2}}{q_{1}}R^{3}(r_{h}).$ (62)
Setting $q_{1}=1$, we obtain:
$\sigma_{xx}^{(33)}=1-q_{2}R^{3}(r_{h}).$ (63)
$R(r_{h})$ is Ricci scalar on event horizon.
$R(r_{h})=-\frac{\Psi(r_{h})^{3}}{r_{h}^{6}},$ (64)
so the color non-abelian DC conductivity is as,
$\sigma_{xx}^{(33)}=1$-
$q_{2}\bigg[\frac{(r_{h}\left(f^{\prime}(r_{h})(4-3r_{h}H^{\prime}(r_{h}))+r_{h}f^{\prime\prime}(r_{h})\right)+2f\left((-1+r_{h}H^{\prime}(r_{h}))^{2}-r_{h}^{2}H^{\prime\prime}(r_{h})\right)))^{3}}{r_{h}^{6}}\bigg].$ (65)
This result shows that the conductivity bound is violated for the non-abelian non-minimal $R^{3}F^{2}$ black brane theory when $q_{2}R^{3}(r_{h})>0$.
Up to first order of $q_{2}$, the color non-abelian DC conductivity is violated for $q_{2}>0$ as,
$\sigma_{xx}^{(33)}=1-q_{2}\frac{1728}{L^{6}}.$ (66)
In the limit $q_{2}\to 0$, we recover:
$\sigma_{xx}^{(33)}=1.$ (67)
4 Shear Viscosity to Entropy Density Ratio
The ratio of shear viscosity to entropy density $\eta/s$ is a fundamental transport coefficient that characterizes the fluid dynamics of the dual field theory. We compute this ratio using the membrane paradigm approach for our non-minimal $R^{3}F^{2}$ black brane.
We consider metric perturbations of the form:
$g_{\mu\nu}\to g_{\mu\nu}+h_{\mu\nu},$ (68)
with the specific perturbation $h_{xy}=h_{xy}(t,r)=\frac{r^{2}}{L^{2}}\phi(r)e^{-i\omega t}$. We work in the radial gauge where $h_{r\mu}=0$.
The quadratic action for the perturbation $\phi(r)$ can be obtained by expanding the action (1) to second order. For our non-minimal model, the effective action takes the form:
$S^{(2)}=\frac{1}{2\kappa}\int d^{4}x\sqrt{-g}\,\mathcal{K}(r)\left[g^{rr}(\partial_{r}\phi)^{2}+g^{tt}(\partial_{t}\phi)^{2}\right],$ (69)
where $\mathcal{K}(r)$ is the effective coupling that encodes the non-minimal corrections.
For the $R^{3}F^{2}$ coupling, we compute $\mathcal{K}(r)$ by expanding the action to second order in $\phi$ and extracting the coefficient of the kinetic term. The calculation yields:
$\mathcal{K}(r)=1+q_{2}\mathcal{C}(r)+\mathcal{O}(q_{2}^{2}),$ (70)
where the correction term $\mathcal{C}(r)$ is given by:
$\mathcal{C}(r)=-12R(r)^{2}F^{(a)}_{\mu\nu}F^{(a)\mu\nu}-24R(r)^{3}.$ (71)
Using the membrane paradigm, the shear viscosity can be expressed as a horizon quantity:
$\eta=\frac{1}{2\kappa}\left(\sqrt{-g}g^{rr}g^{xx}\mathcal{K}(r)\right)\Bigg|_{r=r_{h}}.$ (72)
Substituting our background metric (8) and evaluating at the horizon:
$\eta$ $=\frac{1}{2\kappa}\left(\frac{r_{h}^{2}}{L^{2}}e^{-H(r_{h})}\mathcal{K}(r_{h})\right)$ (73)
$=\frac{1}{2\kappa}\left(\frac{r_{h}^{2}}{L^{2}}\right)\left[1+q_{2}\mathcal{C}(r_{h})+\mathcal{O}(q_{2}^{2})\right],$ (74)
where we used $e^{-H(r_{h})}\approx 1$ to first order in $q_{2}$.
The entropy density $s$ is given by the Bekenstein-Hawking formula:
$s=\frac{1}{4G}\left(\frac{r_{h}^{2}}{L^{2}}\right)=\frac{2\pi}{\kappa}\left(\frac{r_{h}^{2}}{L^{2}}\right).$ (75)
We consider $\kappa=8\pi G$.
Therefore, the ratio $\eta/s$ is:
$\frac{\eta}{s}=\frac{1}{4\pi}\left[1+q_{2}\mathcal{C}(r_{h})+\mathcal{O}(q_{2}^{2})\right].$ (76)
Now we compute $\mathcal{C}(r_{h})$ explicitly using our background solution. At the horizon $r=r_{h}$, we have:
$R(r_{h})$ $=-12/L^{2}+\mathcal{O}(q_{2}),$ (77)
$F^{(a)}_{\mu\nu}F^{(a)\mu\nu}(r_{h})$ $=-2h^{\prime}(r_{h})^{2}e^{2H(r_{h})}+\mathcal{O}(q_{2})=-\frac{2Q^{2}}{r_{h}^{4}}+\mathcal{O}(q_{2}).$ (78)
Substituting these into $\mathcal{C}(r_{h})$:
$\mathcal{C}(r_{h})$ $=-12\left(-\frac{12}{L^{2}}\right)^{2}\left(-\frac{2Q^{2}}{r_{h}^{4}}\right)-24\left(-\frac{12}{L^{2}}\right)^{3}$ (79)
$=-12\left(\frac{144}{L^{4}}\right)\left(-\frac{2Q^{2}}{r_{h}^{4}}\right)-24\left(-\frac{1728}{L^{6}}\right)$ (80)
$=\frac{3456Q^{2}}{L^{4}r_{h}^{4}}+\frac{41472}{L^{6}}.$ (81)
Therefore, the shear viscosity to entropy density ratio to first order in $q_{2}$ is:
$\frac{\eta}{s}=\frac{1}{4\pi}\left[1+q_{2}\left(\frac{3456Q^{2}}{L^{4}r_{h}^{4}}+\frac{41472}{L^{6}}\right)+\mathcal{O}(q_{2}^{2})\right].$ (82)
This result shows that the Kovtun-Son-Starinets (KSS) bound $\eta/s\geq 1/(4\pi)$ is satisfied for $q_{2}>0$. However, for $q_{2}<0$, the bound is violated.
In the limit $q_{2}\to 0$, we recover the universal value for Einstein gravity:
$\frac{\eta}{s}=\frac{1}{4\pi}.$ (83)
5 Conclusion
In this paper, we have constructed a non-minimal $R^{3}F^{(a)}_{\mu\alpha}F^{(a)\mu\alpha}$ black brane solution in AdS spacetime in four dimensions. Due to the complexity of the field equations, we employed a perturbative approach to obtain solutions accurate to first order in the coupling parameter $q_{2}$. Using the framework of gauge/gravity duality, we investigated two important holographic transport coefficients: the color non-abelian DC conductivity and the ratio of shear viscosity to entropy density.
Our calculations demonstrate that the non-minimal $R^{3}F^{2}$ coupling significantly affects these transport coefficients. Specifically, we found that the conductivity bound is violated for positive values of $q_{2}$, while the Kovtun-Son-Starinets bound for shear viscosity is violated for negative $q_{2}$. These violations highlight how higher-order non-minimal couplings between curvature and matter fields can modify the transport properties of the dual field theory.
In the limit $q_{2}\to 0$, our results smoothly reduce to those of the standard Einstein-Yang-Mills theory, with the conductivity bound saturated and the KSS bound maintaining its universal value of $1/4\pi$. This consistency check validates our perturbative approach and demonstrates the robustness of our methodology.
The violations of universal bounds observed in our model suggest that non-minimal $R^{3}F^{2}$ couplings provide a rich framework for exploring beyond-standard-model physics within the holographic context. Future work could extend this analysis to higher-order perturbations or investigate other holographic observables to further elucidate the implications of such non-minimal couplings in the context of gauge/gravity duality.
Data availability statement
All data that support the findings of this study are included within the article.
Appendix A The relation of $\mathcal{L}(t)$
$\mathcal{L}(t)=q_{1}\mathcal{A}(t)+q_{2}\mathcal{B}(t),$ (A-1)
where,
$\mathcal{A}(t)$ $=-t^{4}f^{\prime\prime}(t)-2t^{6}f(t)\left[1+tH^{\prime}(t)\right],$ (A-2)
$\mathcal{B}(t)$ $=\mathcal{B}_{1}(t)+\mathcal{B}_{2}(t)+\mathcal{B}_{3}(t)+\mathcal{B}_{4}(t)+\mathcal{B}_{5}(t),$ (A-3)
with,
$\mathcal{B}_{1}(t)$ $=-4t^{4}f^{\prime\prime}(t)^{4}-4f^{\prime}(t)^{4}(4-3tH^{\prime}(t))^{2}\Theta_{1}(t),$
$\mathcal{B}_{2}(t)$ $=4tf^{\prime}(t)^{3}(-4+3tH^{\prime}(t))\left[f^{\prime\prime}(t)\Theta_{2}(t)+3t(4-3tH^{\prime}(t))f^{\prime\prime\prime}(t)\right],$
$\mathcal{B}_{3}(t)$ $=12t^{2}f^{\prime}(t)^{2}f^{\prime\prime}(t)\left[f^{\prime\prime}(t)\Theta_{3}(t)+2t(-4+3tH^{\prime}(t))f^{\prime\prime\prime}(t)\right],$
$\mathcal{B}_{4}(t)$ $=t^{3}f^{\prime}(t)\left[2q_{1}+tH^{\prime}(t)(q_{1}+76q_{2}f^{\prime\prime}(t)^{3})-4q_{2}f^{\prime\prime}(t)^{2}(22f^{\prime\prime}(t)+3tf^{\prime\prime\prime}(t))\right],$
$\mathcal{B}_{5}(t)$ $=-64q_{2}f(t)^{4}\Gamma(t)^{2}\Theta_{4}(t)$
$\quad+32q_{2}f(t)^{3}t\Gamma(t)\left\{-t\left[3t\Gamma(t)f^{\prime\prime\prime}(t)+f^{\prime\prime}(t)\Theta_{5}(t)\right]+f^{\prime}(t)\Theta_{6}(t)\right\}$
$\quad-2f(t)t^{3}\left[q_{1}t^{3}(1+tH^{\prime}(t))-4q_{2}\left(f^{\prime}(t)(-4+3tH^{\prime}(t))-tf^{\prime\prime}(t)\right)\right]\times$
$\quad\left[t^{2}f^{\prime\prime}(t)\left(f^{\prime\prime}(t)\Theta_{7}(t)-3tf^{\prime\prime\prime}(t)\right)+f^{\prime}(t)^{2}\Theta_{8}(t)-tf^{\prime}(t)\Theta_{9}(t)\right]$
$\quad-48q_{2}f(t)^{2}t^{2}\left[t^{2}f^{\prime\prime}(t)\Theta_{10}(t)+f^{\prime}(t)^{2}\Theta_{11}(t)+tf^{\prime}(t)\Theta_{12}(t)\right],$
and
$\Gamma(t)$ $=(1-tH^{\prime}(t))^{2}-t^{2}H^{\prime\prime}(t),$
$\Theta_{1}(t)$ $=-14+t\left[H^{\prime}(t)(-10+9tH^{\prime}(t))-15tH^{\prime\prime}(t)\right],$
$\Theta_{2}(t)$ $=28+57tH^{\prime}(t)(-2+tH^{\prime}(t))-30t^{2}H^{\prime\prime}(t),$
$\Theta_{3}(t)$ $=-38+t\left[74H^{\prime}(t)-32tH^{\prime}(t)^{2}+5tH^{\prime\prime}(t)\right],$
$\Theta_{4}(t)$ $=7+t\left[-tH^{\prime}(t)^{2}+t^{2}H^{\prime}(t)^{3}-7H^{\prime}(t)(1+t^{2}H^{\prime\prime}(t))+t(5H^{\prime\prime}(t)+3tH^{\prime\prime\prime}(t))\right],$
$\Theta_{5}(t)$ $=4+t\left[8t^{2}H^{\prime}(t)^{3}+t^{3}H^{\prime}(t)^{4}+H^{\prime}(t)(14-20t^{2}H^{\prime\prime}(t))\right$.
$\quad\left.-tH^{\prime}(t)^{2}(27+2t^{2}H^{\prime\prime}(t))+t\left(H^{\prime\prime}(t)(19+t^{2}H^{\prime\prime}(t))+6tH^{\prime\prime\prime}(t)\right)\right],$
$\Theta_{6}(t)$ $=-58+t\left[13t^{3}H^{\prime}(t)^{4}+t^{4}H^{\prime}(t)^{5}-t^{2}H^{\prime}(t)^{3}(53+8t^{2}H^{\prime\prime}(t))\right$.
$\quad+tH^{\prime}(t)^{2}(17-48t^{2}H^{\prime\prime}(t)+3t^{3}H^{\prime\prime\prime}(t))$
$\quad+t\left(H^{\prime\prime}(t)(-49+11t^{2}H^{\prime\prime}(t))-3t(7+t^{2}H^{\prime\prime}(t))H^{\prime\prime\prime}(t)\right)$
$\quad\left.+H^{\prime}(t)\left(80+t^{2}\left(H^{\prime\prime}(t)(123+7t^{2}H^{\prime\prime}(t))+12tH^{\prime\prime\prime}(t)\right)\right)\right],$
$\Theta_{7}(t)$ $=-8+t\left[H^{\prime}(t)(4+3tH^{\prime}(t))-3tH^{\prime\prime}(t)\right],$
$\Theta_{8}(t)$ $=-20+t\left[-33t^{2}H^{\prime}(t)^{3}+21t^{3}H^{\prime}(t)^{4}-3tH^{\prime}(t)^{2}(24+23t^{2}H^{\prime\prime}(t))\right$.
$\quad+6t\left(H^{\prime\prime}(t)(7+5t^{2}H^{\prime\prime}(t))-2tH^{\prime\prime\prime}(t)\right)$
$\quad\left.+H^{\prime}(t)\left(118+75t^{2}H^{\prime\prime}(t)+9t^{3}H^{\prime\prime\prime}(t)\right)\right],$
$\Theta_{9}(t)$ $=10+t\left[-51tH^{\prime}(t)^{2}+30t^{2}H^{\prime}(t)^{3}-2H^{\prime}(t)(1+18t^{2}H^{\prime\prime}(t))\right$.
$\quad\left.+3t\left(7H^{\prime\prime}(t)+tH^{\prime\prime\prime}(t)\right)\right],$
$\Theta_{10}(t)$ $=-4+t\left[3t^{2}H^{\prime}(t)^{3}+t^{3}H^{\prime}(t)^{4}-5H^{\prime}(t)(-3+t^{2}H^{\prime\prime}(t))\right$.
$\quad\left.-tH^{\prime}(t)^{2}(15+2t^{2}H^{\prime\prime}(t))+t\left(H^{\prime\prime}(t)(7+t^{2}H^{\prime\prime}(t))+tH^{\prime\prime\prime}(t)\right)\right],$
$\Theta_{11}(t)$ $=38+t\left[-t^{4}H^{\prime}(t)^{5}+5t^{5}H^{\prime}(t)^{6}-27t^{3}H^{\prime}(t)^{4}(3+t^{2}H^{\prime\prime}(t))\right$.
$\quad+t^{2}H^{\prime}(t)^{3}(185+27t^{2}H^{\prime\prime}(t)+6t^{3}H^{\prime\prime\prime}(t))$
$\quad+t\left(H^{\prime\prime}(t)(55-t^{2}H^{\prime\prime}(t)(24+5t^{2}H^{\prime\prime}(t)))+8t(1+t^{2}H^{\prime\prime}(t))H^{\prime\prime\prime}(t)\right)$
$\quad+tH^{\prime}(t)^{2}\left(-112+t^{2}\left(H^{\prime\prime}(t)(121+27t^{2}H^{\prime\prime}(t))-11tH^{\prime\prime\prime}(t)\right)\right)$
$\quad\left.-2H^{\prime}(t)\left(17+t^{2}\left(H^{\prime\prime}(t)(93+10t^{2}H^{\prime\prime}(t))+t(1+3t^{2}H^{\prime\prime}(t))H^{\prime\prime\prime}(t)\right)\right)\right],$
$\Theta_{12}(t)$ $=-2+t\left[6t^{3}H^{\prime}(t)^{4}-7t^{4}H^{\prime}(t)^{5}+2t^{2}H^{\prime}(t)^{3}(26+9t^{2}H^{\prime\prime}(t))\right$.
$\quad+2t\left(17H^{\prime\prime}(t)+t(3+t^{2}H^{\prime\prime}(t))H^{\prime\prime\prime}(t)\right)$
$\quad-2tH^{\prime}(t)^{2}\left(50+t^{2}\left(H^{\prime\prime}(t)+tH^{\prime\prime\prime}(t)\right)\right)$
$\quad\left.-H^{\prime}(t)\left(-51+t^{2}\left(H^{\prime\prime}(t)(60+11t^{2}H^{\prime\prime}(t))+2tH^{\prime\prime\prime}(t)\right)\right)\right].$